# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ C.3.2. Proof of Lemma 6 [Pruned possible graphs for $C^{-1}$]
Next, assume that $H^{-1}$ is the graph in either Figure 19(f), 19(g), 19(h), or 19(i), or $H^{-1}$ is the graph in 19(l) and $uw_{2}$ is an edge. Since $I^{-1}$ has an $f$-preserving forbidden minor, $G\cup f$ has an $f$-preserving forbidden minor in which some edge $e$ with $v$ as an endpoint is contracted. We will show that $e$ is reducing. The previous observation shows that $e$ is not of Type (1). Let $[G\cup f]$ be the minor obtained by contracting $e$. Note that $f$ is retained in $[G\cup f]$, and so $e$ is not of Type (3). We complete the proof by showing that $e$ is not of Type (2). Since $G^{\prime}$ is the union of $C^{-1}$ and some non-empty subset of connected components of $(G\cup f)\setminus C^{-1}$ and since $G\cup f$ is an atom, Lemma 61 in Appendix E states that, for any cms $E$ of $G^{\prime}$, each connected component of $G^{\prime}\setminus E$ contains some vertex of $C^{-1}\setminus E$. Along with the observation that $C^{-1}$ is an atom, this shows that $G^{\prime}$ is an atom. Since $G^{\prime}$ is weakly retained in $[G\cup f]$, we get that $[G^{\prime}]$ is an atom. Additionally, note that $[H^{-1}]$ is an atom. Since $[C^{-1}]$ is not a clique, $[G\cup f]=[G^{\prime}]\cup[H^{-1}]$, and $[C^{-1}]=[G^{\prime}]\cap[H^{-1}]$, Lemma 64 shows that $[G\cup f]$ is an atom, and so $e$ is not of Type (2). Therefore, $e$ is reducing, as desired.
Finally, assume that $H^{-1}$ is the graph in Figure 19(j). If either $uw_{1}$ or $vw_{3}$ is not of Type (1), then an argument similar to those above shows that at least one of these edges is reducing. Hence, we complete the proof by showing that at least one of these edges is not of Type (1). By definition, $I^{-1}$ has an $f$-preserving forbidden minor $[I^{-1}]_{A}$. If either $w_{1}w_{4}$ or $w_{3}w_{5}$ is preserved in $[I^{-1}]_{A}$, then there exists an extension $[G\cup f]$ of $[I^{-1}]_{A}$ in which either $uw_{1}$ or $vw_{3}$ is contracted, and we are done. Otherwise, note that $w_{2}w_{3}$ is preserved in $[I^{-1}]_{A}$ since $f$ is preserved in $[I^{-1}]_{A}$. Therefore, there exists an extension $[G\cup f]$ of $[I^{-1}]_{A}$ in which both $uw_{1}$ and $vw_{3}$ are contracted, and hence neither of these edges is of Type (1). Similar arguments apply if either $H^{-1}$ is the graph in Figure 19(k) or $H^{-1}$ is the graph in Figure 19(l) and $uw_{2}$ is a nonedge.
@@
C.3.3. Proof of Lemma 7 [Unpruned possible graphs for top-level $H^{-1}$]
Here we prove Lemma 7 from Section 4. Consider a minimal graph-nonedge pair $(G,f)$ with an expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$. This lemma enumerates all possible graphs for $H^{-1}$. Its proof is highly technical and has many cases, which are split amongst Lemmas 39, 40, and 41, below. The proofs of these lemmas require Lemmas 36 and 38, below, which are proved using Lemmas 35 and 37, also below.
@@ C.3.3. Proof of Lemma 7 [Unpruned possible graphs for top-level $H^{-1}$]
Assume to the contrary that $F$ is a non-empty set of nonedges of $C^{-1}$ that satisfy Statements (i)-(iii) but $I^{-1}\setminus C^{-1}$ contains more than one vertex. Then, there exists some edge $e$ in $I^{-1}\setminus C^{-1}$. We will show that $e$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Since neither endpoint of $e$ is contained in $H^{-1}$, $f$ is retained in be the minor $[G\cup f]_{A}$ obtained by contracting $e$, and so $e$ is not of Type (3). Combining this with Statements (i)-(iii) shows that $[G\cup f]_{A}$ has an $f$-preserving forbidden minor, and so $e$ is not of Type (1). We complete the proof by showing that $e$ is not of Type (2).
Since $G\cup f$ is an atom, any cms of $[G\cup f]_{A}$ must contain the vertex $[e]_{A}$. Hence, since $[C^{-1}]_{A}$ is a separator of $[G\cup f]_{A}$, every cms of $[G\cup f]_{A}$ is contained in $[I^{-1}]_{A}$. Furthermore, observe that $G^{\prime}=(G\cup f)\setminus(I^{-1}\setminus C^{-1})$ is weakly retained in $[G\cup f]_{A}$. Therefore, Lemma 31 applies and shows that $[G^{\prime}]_{A}$ is contained in some atom $K$ of $[G\cup f]_{A}$. Note that $K$ contains $[H^{-1}]_{A}$, and so $K\cup[F]_{A}$ has an $[f]_{A}$-preserving forbidden minor, by Statement (iii). Also, since any cms of $[G\cup f]_{A}$ contains $[e]_{A}$, $K$ contains $[e]_{A}$. For each endpoint $y$ of each nonedge in $[F]_{A}$, Statement (ii) can be used to obtain some path in $[G\cup f]_{A}$ between $[e]_{A}$ and $y$ that contains exactly one vertex in $[H^{-1}]_{A}$. Consequently, Lemma 66 in Section E allows us to choose this path to be contained in $K$. Combining this with Statements (i)-(iii) shows that $K$ has an $[f]_{A}$-preserving forbidden minor, and thus $e$ is not of Type (2). ∎
@@
  1. (1)
    If $C^{-1}$ is the graph in either Figure 13(a), 13(c), or 13(d), then, for any non-empty subset $F$ of nonedges of $C^{-1}$, $H^{-1}\cup F$ has no $f$-preserving forbidden minor.
  2. (2)
@@
Let $I^{-1}$ be the expanded minor component of $C^{-1}$ and consider the minor $[G\cup f]$ obtained by contracting every edge in $I^{-1}\setminus C^{-1}$. Also, let $[G\cup f]_{A}$ be an $f$-separating minor such that $C$ is one of its $f$-separating cms’s. If Statement (1) is false, then some non-empty set $F$ of nonedges $C^{-1}$ is such that $H^{-1}\cup F$ has an $f$-preserving forbidden minor. Using Lemma 4 Statement (2), we see that the vertex $[I^{-1}\setminus C^{-1}]$ neighbors the endpoints of each nonedge in $F$. If Statement (2) is false, then, for any nonedge $f^{\prime}$ of $C^{-1}$, $H^{-1}\cup f^{\prime}$ has an $f$-preserving forbidden minor. Using Statement (2) of Lemma 4, we see that the vertex $[I^{-1}\setminus C^{-1}]$ neighbors the endpoints of some nonedge of $C^{-1}$. In both cases, Lemma 35 applies and shows that $I^{-1}\setminus C^{-1}$ contains exactly one vertex. However, since $|V(C^{-1})|\leq 4$, this implies that $[I^{-1}]_{A}$ has no forbidden minor, which is a contradiction. This completes the proof. ∎
@@
First, letting $X=E^{-1}\cap C^{-1}$, we show that $C^{-1}\setminus X$ is non-empty. Since $G\cup f$ is an atom, $E^{-1}$ must contain $ab$. Letting $A=E^{-1}\cap G^{\prime}$, the above facts show that $E^{-1}=A\cup a$. Since $G^{\prime}$ is weakly retained in $[G\cup f]$, $A$ is a clique. Consequently, $X$ is a clique. Also, since $C^{-1}$ is contained in $G^{\prime}$ and is not a clique, by Lemma 6, $C^{-1}\setminus X$ is non-empty.
@@
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component be $H^{-1}$. If $C^{-1}$ is the graph in Figure 13(a) and contains exactly one endpoint of $f$, then $H^{-1}$ is one of the graphs in Figures 15(a), 15(e), or 15(n) such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
@@
Let $f=uv$, where $u$ is not contained in $C^{-1}$. If $u$ is the only vertex of $H^{-1}\setminus C^{-1}$, then the fact that $G\cup f$ is an atom implies that $H^{-1}$ is one of the graphs in Figures 15(a) or 15(n) and has the desired properties. Otherwise, we will consider the two cases for $v$ in $C^{-1}$.
@@
We will show that some $H^{-1}$ is the graph in Figure 15(e) has the desired properties. Wlog, let $v=w_{1}$. By Lemma 5, each vertex of $H^{-1}\setminus(C^{-1}\setminus u)$ neighbors $u$, $w_{1}$, and at least one vertex of $C^{-1}\setminus w_{1}$. First we show that no such vertex neighbors $w_{2}$. If this is not the case, then let $x$ be any such vertex that neighbors $w_{2}$. We will show that $xw_{2}$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{2}$, $xw_{2}$ is not of Type (3). To see that $xw_{2}$ is not of Type (1) or (2), consider the $f$-separating minor $[G\cup f]_{A}$ obtained by contracting some edge in $\{w_{1}w_{2},w_{2}w_{3}\}$ and that has $C$ as one of its $f$-separating cms’s. Since $G\cup f$ is an atom, there exists some path in $G\cup f$ from $u$ to $w_{3}$ that does not contain any vertex in $\{x,w_{1},w_{2}\}$. Also, the definition of $C$ shows that some $C$-component of $[G\cup f]_{A}$ that does not contain $[f]_{A}$ has an $[f]_{A}$-preserving forbidden minor. By combining these facts, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $xw_{2}$ is not of Type (1). Lastly, observe that $G^{\prime}=(G\cup f)\setminus(H^{-1}\setminus C^{-1})$ is weakly retained in $[G\cup f]$. Furthermore, Statement (i) of Lemma 38 is true if we set $ab=xw_{2}$. Consequently, this lemma shows that $xw_{2}$ is not of Type (2), and so the above facts show that $xw_{2}$ is reducing.
@@
Next, we show that $u$ neighbors $w_{2}$. Assume this is not the case. The above argument shows that the neighborhood of any vertex $x$ in $H^{-1}\setminus(C^{-1}\setminus u)$ is $\{u,w_{1},w_{3}\}$. Since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{3}$, $xw_{3}$ is not of Type (3). Also, since some $C$-component of $[G\cup f]_{A}$ that does not contain $[f]_{A}$ has an $[f]_{A}$-preserving forbidden minor, $[G\cup f]$ clearly has an $[f]$-preserving forbidden minor, and so $xw_{3}$ is not of Type (1). Hence, $xw_{3}$ must be of Type (2). Let $E$ be the subgraph of $[G\cup f]$ induced by $\{[w_{1}],[w_{3}]\}$ and let $I^{-1}$ be the expanded minor component of $C^{-1}$. Observe that $I^{-1}$ is preserved in $[G\cup f]$ and $E$ is a $[uy]$-separator for any vertex $[y]$ of $[I^{-1}]\setminus E$. Since $I^{-1}$ has an $f$-preserving forbidden minor, Lemma 27 can be used to show that there exists an atom $M$ of $[G\cup f]$ that has an $[f]$-preserving forbidden minor and such that $E$ is a $[uy]$-separator for any vertex $[y]$ of $M\setminus E$. Furthermore, since $G\cup f$ is an atom and $E$ is a clique separator of $[G\cup f]$ that contains exactly two vertices, it is easy to see that $E$ is a minimal $[uy]$-separator for any vertex $[y]$ of $M\setminus E$. Therefore, $(E,M)$ is an $f$-separating pair of $[G\cup f]$. However, note that the $[f]$-component of $E$ is $[H^{-1}]\setminus[w_{2}]$, and so the expanded $f$-component of $E^{-1}$ is $H^{-1}\setminus w_{2}$. This shows that $C^{-1}$ is not in the top-level of $G\cup f$, which is a contradiction.
@@
Finally, we show that either $u$ neighbors $w_{3}$ or $H^{-1}\setminus(C^{-1}\setminus u)$ contains at least two vertices, which completes the proof. If this is not the case, then $H^{-1}\setminus(C^{-1}\setminus u)$ contains exactly one vertex $x$ and all paths in $H^{-1}$ between $u$ and $w_{3}$ contain some vertex in $\{x,w_{1},w_{2}\}$. If $w_{1}w_{2}$ is contracted in $[G\cup f]_{A}$, then $C^{-1}$ is clearly not in the top-level of $G\cup f$, which is a contradiction. Otherwise, we arrive at a similar contradiction by considering the minor $[G\cup f]$ obtained by contracting $xw_{1}$ and the subgraph of $[G\cup f]$ induced by the set $\{[w_{1}],[w_{2}]\}$.
@@
Wlog, assume that $v$ neighbors $w_{2}$. Since $G\cup f$ is an atom, $u$ neighbors some vertex in $\{w_{1},w_{3}\}$, wlog say $w_{1}$, and either $u$ or $v$ neighbors $w_{3}$. Consider the case where $v$ neighbors $w_{3}$. If $v$ neighbors $w_{1}$, then $u$ must neighbor $w_{3}$. If, additionally, $u$ neighbors $w_{2}$, then $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 36. Otherwise, an argument as in Case 1 of Lemma 39 shows that $vw_{2}$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Therefore, assume $v$ does not neighbor $w_{1}$. If, additionally, $u$ does not neighbor $w_{3}$, then a similar argument shows that $uw_{1}$ is reducing, which is a contradiction. Otherwise, if $u$ neighbors $w_{2}$, then $\{u,w_{2},w_{3}\}$ is a cs of $G\cup f$, which is a contradiction. Else, a similar argument shows that $vw_{2}$ is reducing, which is a contradiction.
Next, consider the case where $u$ neighbors $w_{3}$ and $v$ does not. If $u$ neighbors $w_{2}$, then the $\{u,w_{1},w_{2}\}$ is a cs of $G\cup f$, which is a contradiction. Otherwise, a similar argument shows that $vw_{2}$ is reducing, which is a contradiction.
@@
Since $G\cup f$ is an atom, if either $u$ or $v$ neighbors both $w_{1}$ and $w_{3}$, then $u$ and $v$ each neighbor both $w_{1}$ and $w_{3}$. Consider the $f$-separating minor $[G\cup f]_{A}$ obtained by contracting some edge in $\{w_{1}w_{2},w_{2}w_{3}\}$ and that has $C$ as one of its $f$-separating cms’s. The definition of $C$ shows that some $C$-component of $[G\cup f]_{A}$ that does not contain $[f]_{A}$ has an $[f]_{A}$-preserving forbidden minor. This implies that some edge $e\in\{vw_{1},vw_{3}\}$ is contracted in $[G\cup f]_{A}$, and so $e$ is not of Type (1). Note that the unique atom of the minor $[G\cup f]$, obtained by contracting $e$, that contains $[u]$ is $K_{3}$, and hence $e$ is clearly not of Type (3). Consequently, $e$ is of Type (2). Let $E$ be the cms of $[G\cup f]$ induced by the set $\{[w_{1}],[w_{3}]\}$. Using an argument as in Case 1 of Lemma 39, we see that $E^{-1}$ is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$. However, this contradicts the fact that $C^{-1}$ is in the top-level of $G\cup f$.

### v2

@@ B.2   Proofs of lemmas for Proposition 8
Next, assume that $H^{-1}$ is the graph in either Figure 24(f), 24(g), 24(h), or 24(i), or $H^{-1}$ is the graph in 24(l) and $uw_{2}$ is an edge. Since $I^{-1}$ has an $f$-preserving forbidden minor, $G\cup f$ has an $f$-preserving forbidden minor in which some edge $e$ with $v$ as an endpoint is contracted. We will show that $e$ is reducing. The previous observation shows that $e$ is not of Type (1). Let $[G\cup f]$ be the minor obtained by contracting $e$. Note that $f$ is retained in $[G\cup f]$, and so $e$ is not of Type (3). We complete the proof by showing that $e$ is not of Type (2). Since $G^{\prime}$ is the union of $C^{-1}$ and some non-empty subset of connected components of $(G\cup f)\setminus C^{-1}$ and since $G\cup f$ is an atom, Lemma 63 in C states that, for any CMS $E$ of $G^{\prime}$, each connected component of $G^{\prime}\setminus E$ contains some vertex of $C^{-1}\setminus E$. Along with the observation that $C^{-1}$ is an atom, this shows that $G^{\prime}$ is an atom. Since $G^{\prime}$ is weakly retained in $[G\cup f]$, we get that $[G^{\prime}]$ is an atom. Additionally, note that $[H^{-1}]$ is an atom. Since $[C^{-1}]$ is not a clique, $[G\cup f]=[G^{\prime}]\cup[H^{-1}]$, and $[C^{-1}]=[G^{\prime}]\cap[H^{-1}]$, Lemma 66 in C shows that $[G\cup f]$ is an atom, and so $e$ is not of Type (2). Therefore, $e$ is reducing, as desired.
Finally, assume that $H^{-1}$ is the graph in Figure 24(j). If either $uw_{1}$ or $vw_{3}$ is not of Type (1), then an argument similar to those above shows that at least one of these edges is reducing. Hence, we complete the proof by showing that at least one of these edges is not of Type (1). By definition, $I^{-1}$ has an $f$-preserving forbidden minor $[I^{-1}]_{1}$. If either $w_{1}w_{4}$ or $w_{3}w_{5}$ is preserved in $[I^{-1}]_{1}$, then there exists an extension $[G\cup f]$ of $[I^{-1}]_{1}$ in which either $uw_{1}$ or $vw_{3}$ is contracted, and we are done. Otherwise, note that $w_{2}w_{3}$ is preserved in $[I^{-1}]_{1}$ since $f$ is preserved in $[I^{-1}]_{1}$. Therefore, there exists an extension $[G\cup f]$ of $[I^{-1}]_{1}$ in which both $uw_{1}$ and $vw_{3}$ are contracted, and hence neither of these edges is of Type (1). Similar arguments apply if either $H^{-1}$ is the graph in Figure 24(k) or $H^{-1}$ is the graph in Figure 24(l) and $uw_{2}$ is a nonedge.
@@
The next subsection is dedicated to proving Lemma 31 from Section 5.1.
B.2.1   Proof of Lemma 31 [Unpruned possible graphs for top-level $H^{-1}$]
The proof is highly technical and has many cases, which are split among Lemmas 49, 50, and 51, below. The proofs of these lemmas require Lemmas 46 and 48, below. Lemma 46 is proved using Lemma 45, also below.
@@ B.2.1   Proof of Lemma 31 [Unpruned possible graphs for top-level $H^{-1}$]
Assume to the contrary that $F$ is a non-empty set of nonedges of $C^{-1}$ that satisfy Statements (i)-(iii) but $I^{-1}\setminus C^{-1}$ contains more than one vertex. Then, there exists some edge $e$ in $I^{-1}\setminus C^{-1}$. We will show that $e$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Since neither endpoint of $e$ is contained in $H^{-1}$, $f$ is retained in be the minor $[G\cup f]_{1}$ obtained by contracting $e$, and so $e$ is not of Type (3). Combining this with Statements (i)-(iii) shows that $[G\cup f]_{1}$ has an $f$-preserving forbidden minor, and so $e$ is not of Type (1). We complete the proof by showing that $e$ is not of Type (2).
Since $G\cup f$ is an atom, any CMS of $[G\cup f]_{1}$ must contain the vertex $[e]_{1}$. Hence, since $[C^{-1}]_{1}$ is a separator of $[G\cup f]_{1}$, every CMS of $[G\cup f]_{1}$ is contained in $[I^{-1}]_{1}$. Furthermore, observe that $G^{\prime}=(G\cup f)\setminus(I^{-1}\setminus C^{-1})$ is weakly retained in $[G\cup f]_{1}$. Therefore, Lemma 43 applies and shows that $[G^{\prime}]_{1}$ is contained in some atom $K$ of $[G\cup f]_{1}$. Note that $K$ contains $[H^{-1}]_{1}$, and so $K\cup[F]_{1}$ has an $[f]_{1}$-preserving forbidden minor, by Statement (iii). Also, since any CMS of $[G\cup f]_{1}$ contains $[e]_{1}$, $K$ contains $[e]_{1}$. For each endpoint $y$ of each nonedge in $[F]_{1}$, Statement (ii) can be used to obtain some path in $[G\cup f]_{1}$ between $[e]_{1}$ and $y$ that contains exactly one vertex in $[H^{-1}]_{1}$. Consequently, Lemma 68 in Section C allows us to choose this path to be contained in $K$. Combining this with Statements (i)-(iii) shows that $K$ has an $[f]_{1}$-preserving forbidden minor, and thus $e$ is not of Type (2). ∎
@@
  1. 1.
    If $C^{-1}$ is the graph in either Figure 22(a), 22(c), or 22(d), then, for any non-empty subset $F$ of nonedges of $C^{-1}$, $H^{-1}\cup F$ has no $f$-preserving forbidden minor.
  2. 2.
@@
Let $I^{-1}$ be the expanded minor component of $C^{-1}$ and consider the minor $[G\cup f]$ obtained by contracting every edge in $I^{-1}\setminus C^{-1}$. Also, let $[G\cup f]_{1}$ be an $f$-separating minor such that $C$ is one of its $f$-separating CMSs. If Statement (1) is false, then some non-empty set $F$ of nonedges $C^{-1}$ is such that $H^{-1}\cup F$ has an $f$-preserving forbidden minor. Using Lemma 25 Statement (2), we see that the vertex $[I^{-1}\setminus C^{-1}]$ neighbors the endpoints of each nonedge in $F$. If Statement (2) is false, then, for any nonedge $f^{\prime}$ of $C^{-1}$, $H^{-1}\cup f^{\prime}$ has an $f$-preserving forbidden minor. Using Statement (2) of Lemma 25, we see that the vertex $[I^{-1}\setminus C^{-1}]$ neighbors the endpoints of some nonedge of $C^{-1}$. In both cases, Lemma 45 applies and shows that $I^{-1}\setminus C^{-1}$ contains exactly one vertex. However, since $|V(C^{-1})|\leq 4$, this implies that $[I^{-1}]_{1}$ has no forbidden minor, which is a contradiction. This completes the proof. ∎
Lemma 48 is proved using Lemma 47, below.
@@
First, letting $X=E^{-1}\cap C^{-1}$, we show that $C^{-1}\setminus X$ is non-empty. Since $G\cup f$ is an atom, $E^{-1}$ must contain $ab$. Let $A=E^{-1}\cap G^{\prime}$. The above facts show that $E^{-1}=A\cup a$. Since $G^{\prime}$ is weakly retained in $[G\cup f]$, $A$ is a clique. Consequently, $X$ is a clique. Also, since $C^{-1}$ is contained in $G^{\prime}$ and is not a clique, by Lemma 30, $C^{-1}\setminus X$ is non-empty.
@@
Let $C^{-1}$ be a top-level expanded $f$-separating CMS of a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component be $H^{-1}$. If $C^{-1}$ is the graph in Figure 22(a) and contains exactly one endpoint of $f$, then $H^{-1}$ is one of the graphs in Figures 20(a), 20(e), or 20(n) with $f$ as the red edge and $C^{-1}$ as the subgraph induced by $\{w_{i}\}$.
@@
Let $f=uv$, where $u$ is not contained in $C^{-1}$. If $u$ is the only vertex of $H^{-1}\setminus C^{-1}$, then the fact that $G\cup f$ is an atom implies that $H^{-1}$ is one of the graphs in Figures 20(a) or 20(n) and has the desired properties. Otherwise, we will consider the two cases for $v$ in $C^{-1}$.
@@
We will show that some $H^{-1}$ is the graph in Figure 20(e) has the desired properties. Wlog, let $v=w_{1}$. By Lemma 28, each vertex of $H^{-1}\setminus(C^{-1}\setminus u)$ neighbors $u$, $w_{1}$, and at least one vertex of $C^{-1}\setminus w_{1}$. First we show that no such vertex neighbors $w_{2}$. If this is not the case, then let $x$ be any such vertex that neighbors $w_{2}$. We will show that $xw_{2}$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{2}$, $xw_{2}$ is not of Type (3). To see that $xw_{2}$ is not of Type (1) or (2), consider the $f$-separating minor $[G\cup f]_{1}$ obtained by contracting some edge in $\{w_{1}w_{2},w_{2}w_{3}\}$ and that has $C$ as one of its $f$-separating CMSs. Since $G\cup f$ is an atom, there exists some path in $G\cup f$ from $u$ to $w_{3}$ that does not contain any vertex in $\{x,w_{1},w_{2}\}$. Also, the definition of $C$ shows that some $C$-component of $[G\cup f]_{1}$ that does not contain $[f]_{1}$ has an $[f]_{1}$-preserving forbidden minor. By combining these facts, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $xw_{2}$ is not of Type (1). Lastly, observe that $G^{\prime}=(G\cup f)\setminus(H^{-1}\setminus C^{-1})$ is weakly retained in $[G\cup f]$. Furthermore, Statement (i) of Lemma 48 is true if we set $ab=xw_{2}$. Consequently, this lemma shows that $xw_{2}$ is not of Type (2), and so the above facts show that $xw_{2}$ is reducing.
@@
Next, we show that $u$ neighbors $w_{2}$. Assume this is not the case. The above argument shows that the neighborhood of any vertex $x$ in $H^{-1}\setminus(C^{-1}\setminus u)$ is $\{u,w_{1},w_{3}\}$. Since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{3}$, $xw_{3}$ is not of Type (3). Also, since some $C$-component of $[G\cup f]_{1}$ that does not contain $[f]_{1}$ has an $[f]_{1}$-preserving forbidden minor, $[G\cup f]$ clearly has an $[f]$-preserving forbidden minor, and so $xw_{3}$ is not of Type (1). Hence, $xw_{3}$ must be of Type (2). Let $E$ be the subgraph of $[G\cup f]$ induced by $\{[w_{1}],[w_{3}]\}$ and let $I^{-1}$ be the expanded minor component of $C^{-1}$. Observe that $I^{-1}$ is preserved in $[G\cup f]$ and $E$ is a $[uy]$-separator for any vertex $[y]$ of $[I^{-1}]\setminus E$. Since $I^{-1}$ has an $f$-preserving forbidden minor, Lemma 22 can be used to show that there exists an atom $M$ of $[G\cup f]$ that has an $[f]$-preserving forbidden minor and such that $E$ is a $[uy]$-separator for any vertex $[y]$ of $M\setminus E$. Furthermore, since $G\cup f$ is an atom and $E$ is a clique separator of $[G\cup f]$ that contains exactly two vertices, it is easy to see that $E$ is a minimal $[uy]$-separator for any vertex $[y]$ of $M\setminus E$. Therefore, $(E,M)$ is an $f$-separating pair of $[G\cup f]$. However, note that the $[f]$-component of $E$ is $[H^{-1}]\setminus[w_{2}]$, and so the expanded $f$-component of $E^{-1}$ is $H^{-1}\setminus w_{2}$. This shows that $C^{-1}$ is not in the top-level of $G\cup f$, which is a contradiction.
@@
Finally, we show that either $u$ neighbors $w_{3}$ or $H^{-1}\setminus(C^{-1}\setminus u)$ contains at least two vertices, which completes the proof. If this is not the case, then $H^{-1}\setminus(C^{-1}\setminus u)$ contains exactly one vertex $x$ and all paths in $H^{-1}$ between $u$ and $w_{3}$ contain some vertex in $\{x,w_{1},w_{2}\}$. If $w_{1}w_{2}$ is contracted in $[G\cup f]_{1}$, then $C^{-1}$ is clearly not in the top-level of $G\cup f$, which is a contradiction. Otherwise, we arrive at a similar contradiction by considering the minor $[G\cup f]$ obtained by contracting $xw_{1}$ and the subgraph of $[G\cup f]$ induced by the set $\{[w_{1}],[w_{2}]\}$.
@@
Wlog, assume that $v$ neighbors $w_{2}$. Since $G\cup f$ is an atom, $u$ neighbors some vertex in $\{w_{1},w_{3}\}$, wlog say $w_{1}$, and either $u$ or $v$ neighbors $w_{3}$. Consider the case where $v$ neighbors $w_{3}$. If $v$ neighbors $w_{1}$, then $u$ must neighbor $w_{3}$. If, additionally, $u$ neighbors $w_{2}$, then $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 46. Otherwise, an argument as in Case 1 of Lemma 49 shows that $vw_{2}$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Therefore, assume $v$ does not neighbor $w_{1}$. If, additionally, $u$ does not neighbor $w_{3}$, then a similar argument shows that $uw_{1}$ is reducing, which is a contradiction. Otherwise, if $u$ neighbors $w_{2}$, then $\{u,w_{2},w_{3}\}$ is a clique separator of $G\cup f$, which is a contradiction. Else, a similar argument shows that $vw_{2}$ is reducing, which is a contradiction.
Next, consider the case where $u$ neighbors $w_{3}$ and $v$ does not. If $u$ neighbors $w_{2}$, then the $\{u,w_{1},w_{2}\}$ is a clique separator of $G\cup f$, which is a contradiction. Otherwise, a similar argument shows that $vw_{2}$ is reducing, which is a contradiction.
@@
Since $G\cup f$ is an atom, if either $u$ or $v$ neighbors both $w_{1}$ and $w_{3}$, then $u$ and $v$ each neighbor both $w_{1}$ and $w_{3}$. Consider the $f$-separating minor $[G\cup f]_{1}$ obtained by contracting some edge in $\{w_{1}w_{2},w_{2}w_{3}\}$ and that has $C$ as one of its $f$-separating CMSs. The definition of $C$ shows that some $C$-component of $[G\cup f]_{1}$ that does not contain $[f]_{1}$ has an $[f]_{1}$-preserving forbidden minor. This implies that some edge $e\in\{vw_{1},vw_{3}\}$ is contracted in $[G\cup f]_{1}$, and so $e$ is not of Type (1). Note that the unique atom of the minor $[G\cup f]$, obtained by contracting $e$, that contains $[u]$ is $K_{3}$, and hence $e$ is clearly not of Type (3). Consequently, $e$ is of Type (2). Let $E$ be the CMS of $[G\cup f]$ induced by the set $\{[w_{1}],[w_{3}]\}$. Using an argument as in Case 1 of Lemma 49, we see that $E^{-1}$ is an expanded $f$-separating CMS of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$. However, this contradicts the fact that $C^{-1}$ is in the top-level of $G\cup f$.